### v1

@@ 1 Introduction
In the mid-1990s, Bessis and Zinn-Justin conjectured that the spectrum of the Schrödinger operator with potential $u(x)=ix^{3}$ is real. Later, in the works of Bender and his colleagues, there was started a systematic application to physical problems of differential operators that were ${\mathcal{PT}}$-symmetric, i.e. invariant under simultaneous revearsal of time (${\mathcal{T}}$) and orientation (${\mathcal{P}}$) [2, 3, 4].
@@ 2.3 Theta functions of hyperelliptic Riemann surfaces
$\theta(z)=\sum_{n\in{\mathbb{Z}}^{g}}e^{\pi i\langle Bn,n\rangle+2pii\langle n,z\rangle},$
where $\langle u,v\rangle=\sum_{k+1}^{g}u_{k}v_{k}$ is the standard scalar product of vectors from ${\mathbb{C}}^{g}$.
@@ 2.4.1 The Its–Matveev formula
$U_{k}=\frac{1}{2\pi}\oint\Omega,\ \ \ k=1,\dots,g.$ (5)
@@
$\begin{split}A(P_{1})+\dots+A(P_{g})=-z_{0}-K,\\ K_{j}=\frac{1+B_{jj}}{2}+\sum_{k\neq j}\int_{a_{l}}\left(\omega_{l}(P)A_{j}(P)\right).\end{split}$ (9)
For a generic point $z_{0}\in J(\Gamma)$ these formulas are given by the inversion of the Abel map $A$ from the $g$th symmetric power $\Gamma$ to $J(\Gamma)$:
@@ 3 Spectral curves of Schrödinger ${\mathcal{PT}}$-operators as real Riemann surfaces
1) if all zeros of the polynomial $F(E)$ are real, then the fixed ovals of the antiinvolution $\tau$ divide the surface into two components (the surface is an $M$-curve) and on it one can choose a canonical basis of cycles $a_{1},\dots,b_{g}$ such that
@@ 4 ${\mathcal{PT}}$-potentials
Under the conditions of Propositionl 3 the vector of periods $U$, where $U_{j}=\frac{1}{2\pi}\oint_{b_{j}}\Omega$, $j=1,\dots,g$, is purely imaginary:
@@
We found that under the conditions of Proposition 3, which are satisfied for real and ${\mathcal{PT}}$-potentials, the Itsa–Matveev form takes the form
@@
Let a finite-zone one-dimensional Schrödinger operator be constructed from such spectral data that its spectral curve $\Gamma$ is real (has the form (3), where the polynomial $P_{2g+1}(E)$ has real coefficients) and the divisor of poles $D=P_{1}+\dots+P_{g}$ is such that its image $z_{0}=A(D)$ under the Abel map satisfies (19) (we mean that, according to the proposal 2, a canonical basis of cycles is chosen on $\Gamma$). Then $u(x)$ is a ${\mathcal{PT}}$-potential.
@@ 5 Examples
Case 2. $E_{1}\in{\mathbb{R}},E_{2}=\bar{E}_{3},\mathrm{Im}\,E_{2}\neq 0$. In this case, in the cycle basis sgiven in Proposition 2, the Riemann matrix $B$ has the form
@@
We will not consider two-zone potentials, which could be done by separately distinguishing the case of symmetric curves, for which the theta function is reduced to a function of one variable. But let uss make a few remarks:

### v2

@@ Abstract
Key words: finite-zone Schrödinger operators, PT-symmetry, spectral theory, theta-functional formulas for explicit solutions
2010 Mathematics Subject Classification: 34L40; 47A10; 14K25
Keywords: finite-zone Schrödinger operators, PT-symmetry, spectral theory, theta-functional formulas for explicit solutions
@@ 1 Introduction
In the mid-1990s, Bessis and Zinn-Justin conjectured that the spectrum of the Schrödinger operator with potential $u(x)=ix^{3}$ is real. Later, in the works of Bender and his colleagues, there was started a systematic application to physical problems of differential operators that were ${\mathcal{PT}}$-symmetric, i.e. invariant under simultaneous reversal of time (${\mathcal{T}}$) and orientation (${\mathcal{P}}$) [3, 4, 5].
@@ 2.2 Finite-zone Schrödinger operators
We limited ourselves to the basic facts necessary for considering finite-gap ${\mathcal{PT}}$-potentials. Investigation of one-dimensional Schrödinger operators with complex potentials was started in [16], and finite-zone complex potentials were the first were considered in [2], where the spectral properties of the corresponding operators were studied in detail.
@@ 2.3 Theta functions of hyperelliptic Riemann surfaces
$\theta(z)=\sum_{n\in{\mathbb{Z}}^{g}}e^{\pi i\langle Bn,n\rangle+2\pi i\langle n,z\rangle},$
where $\langle u,v\rangle=\sum_{k=1}^{g}u_{k}v_{k}$ is the standard scalar product of vectors from ${\mathbb{C}}^{g}$.
@@ 2.4.1 The Its–Matveev formula
$U_{k}=\frac{1}{2\pi}\oint_{b_{k}}\Omega,\ \ \ k=1,\dots,g.$ (5)
@@
$\begin{split}A(P_{1})+\dots+A(P_{g})=-z_{0}-K,\\ K_{j}=\frac{1+B_{jj}}{2}+\sum_{l\neq j}\int_{a_{l}}\left(\omega_{l}(P)A_{j}(P)\right).\end{split}$ (9)
For a generic point $z_{0}\in J(\Gamma)$ these formulas are given by the inversion of the Abel map $A$ from the $g$-th symmetric power of $\Gamma$ to $J(\Gamma)$:
@@ 3 Spectral curves of Schrödinger ${\mathcal{PT}}$-operators as real Riemann surfaces
1) if all zeros of the polynomial $P_{2g+1}(E)$ are real, then the fixed ovals of the antiinvolution $\tau$ divide the surface into two components (the surface is an $M$-curve) and on it one can choose a canonical basis of cycles $a_{1},\dots,b_{g}$ such that
@@ 4 ${\mathcal{PT}}$-potentials
Under the conditions of Proposition 3 the vector of periods $U$, where $U_{j}=\frac{1}{2\pi}\oint_{b_{j}}\Omega$, $j=1,\dots,g$, is purely imaginary:
@@
We found that under the conditions of Proposition 3, which are satisfied for real and ${\mathcal{PT}}$-potentials, the Its–Matveev form takes the form
@@
Let a finite-zone one-dimensional Schrödinger operator be constructed from such spectral data that its spectral curve $\Gamma$ is real (has the form (3), where the polynomial $P_{2g+1}(E)$ has real coefficients) and the divisor of poles $D=P_{1}+\dots+P_{g}$ is such that its image $z_{0}=A(D)$ under the Abel map satisfies (19) (we mean that, according to Proposition 2, a canonical basis of cycles is chosen on $\Gamma$). Then $u(x)$ is a ${\mathcal{PT}}$-potential.
@@ 5 Examples
Case 2. $E_{1}\in{\mathbb{R}},E_{2}=\bar{E}_{3},\mathrm{Im}\,E_{2}\neq 0$. In this case, in the cycle basis given in Proposition 2, the Riemann matrix $B$ has the form
@@
We will not consider two-zone potentials, which could be done by separately distinguishing the case of symmetric curves, for which the theta function is reduced to a function of one variable. But let us make a few remarks:
@@ 6 Final remarks
3) The reviewer pointed the author to the article [10], where the problem of the Bloch spectrum for Lamé potentials $u(x)=m(m+1)\omega^{2}\wp(\omega x+z_{0})$, where $\omega$ is the half-period of the Weierstrass function $\wp(z)$, was studied. In this work, the general situation is considered, without additional conditions of reality, and, in particular, for $m=1$, a description of the Bloch spectrum is given using explicit formulas. Although the formula for a finite arc included in the spectrum turns out to be complicated, it significantly simplifies the numerical study.
The author thanks the reviewer who pointed him to the works of [2, 10, 16].